OpenStudy (anonymous):

Solve D =1/4 (e – 3f) for e

OpenStudy (anonymous):

First multiply both sides by 4

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

srry I'm back

OpenStudy (anonymous):

ok...................

OpenStudy (anonymous):

1/4 times 4????

OpenStudy (anonymous):

Nope...... its 1

OpenStudy (anonymous):

wait.................1 times 4?

OpenStudy (anonymous):

D =1/4 (e – 3f) D*4 = 1/4 *4 (e-3f) 4D=(e-3f) 4D+3f=e e=4D+3f

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

Thanks again!!!!

OpenStudy (anonymous):

;)yw